# Shield Synthesis for LTL Modulo Theories

### v1

@@ 3.1 $\textup{LTL}_{\mathcal{T}}$ Shield Construction
Note that if the environment generates an input $v_{\overline{x}}$ (with corresponding $v_{\overline{e}}$) and $D$ responds with an output $v_{\overline{y}}$ such that $\textsf{getchoice}(v_{\overline{y}})=v_{\overline{s}}$, if $W$ permits $(v_{\overline{e}},v_{\overline{s}})$ then the suggested $v_{\overline{y}}$ remains. Otherwise, we use a component find which computes an alternative $v_{\overline{s}}^{\prime}$ such that $(v_{\overline{e}},v_{\overline{s}}^{\prime})$ is in $W$ and a proper value $v_{\overline{y}}$ that corresponds to $v_{\overline{s}}$ is output. Also, note that find can always identify a value $v_{\overline{s}}^{\prime}$ from any $q\in Q_{\textit{now}}$ such that there is a successor $(q^{\prime},v_{\overline{s}}^{\prime})\in\delta(q,v_{\overline{e}})$ by the definition of WR. Given two $v_{\overline{s}}$ and $v_{\overline{s}}^{\prime}$ such that $(v_{\overline{e}},v_{\overline{s}})\in W$ and $(v_{\overline{e}},v_{\overline{s}}^{\prime})\in W$, the criteria (e.g., choice-logic (?)) to choose one is not relevant for this paper and we leave it for future work. Finally, given $v_{\overline{x}}$, the $W$ cannot distinguish two values $v_{\overline{y}}$ and $v_{\overline{y}}^{\prime}$ if $\textsf{getchoice}(v_{\overline{x}},v_{\overline{y}})=\textsf{getchoice}(v_{\overline{x}},v_{\overline{y}}^{\prime})$.
@@
Let $\varphi_{\mathcal{T}}$ be a safety specification and $\varphi_{\mathbb{B}}$ obtained from $\varphi_{\mathcal{T}}$ by Boolean abstraction. Let $S_{\mathcal{T}}$ be synthesized using the winning region $W$ of $\varphi_{\mathbb{B}}$. Then, $D\cdot S_{\mathcal{T}}$ satisfies $\varphi_{\mathcal{T}}$, for any external controller $D$.
@@
Lemma 3 essentially states that a WR controller is less intrusive than one based on a specific Boolean controller $C_{\mathbb{B}}$. However, in practice, a WR is harder to interpret and slower to compute than $C_{\mathbb{B}}$. Note that we presented $S_{\mathcal{T}}$ based on $C_{\mathcal{T}}$ and $S_{\mathcal{T}}$ based on WR, but for classic LTL shielding only the latter case makes sense. This is because in classic LTL using shield $S$ based on $C_{\mathbb{B}}$ is equivalent to forcing $D$ to coincide with $S$ on all time-steps, so $S$ could be directly used ignoring $D$. However, in $\textup{LTL}_{\mathcal{T}}$ there are potentially many different $v_{\overline{y}}$ by $D$ that correspond to a given Boolean $v_{\overline{s}}^{\prime}$ by $C_{\mathbb{B}}$, so even when this architecture forces to mimick $C_{\mathbb{B}}$ in the Boolean domain, it is reasonable to use an external source $D$ of rich value output candidates. Thus, not only we present an extension of shielding to $\textup{LTL}_{\mathcal{T}}$ (with the WR approach), but also this is the first work that is capable to synthetise shields leveraging all the power of classic synthesis (with the $C_{\mathcal{T}}$ approach): for instance, we can use bounded synthesis (?)) instead of necessarily computing a WR.
3.2 Permissive and Intrusive Shields
@@
Computing a Boolean Abstraction boils down to calculating the set VR of valid reactions. In order to speed up the calculation, optimizations are sometimes used (see (?)). For instance, Ex. 1 shows a formula $\varphi_{\mathcal{T}}$ that can be Booleanized into a $\varphi_{\mathbb{B}}$ that contains three environment variables ($e_{0}$, $e_{0}^{+}$ and $e_{1}$) that correspond to the three valid reactions. A closer inspection reveals that the valid reaction associated to $e_{0}^{+}$ (which represents $(x<1)$) is “subsumed” by the reaction associated to $e_{0}$ (which represents $(x<2)$), which is witnessed by $e_{0}$ leaving to the system a strictly smaller set of choices than $e_{0}^{+}$. In the game theoretic sense, an environment that plays $e_{0}^{+}$ is leaving the system strictly more options, so a resulting Booleanizaed formula $\varphi_{\mathbb{B}}^{\prime}$ that simply ignores $e_{0}^{+}$ only limits the environment from playing sub-optimal moves, and $\varphi_{\mathbb{B}}^{\prime}$ is also equi-realizable to $\varphi_{\mathcal{T}}$. From the equi-realizability point of view, one can neglect valid reactions that provide strictly more choices than other valid reactions. However, even though using $\varphi_{\mathbb{B}}^{\prime}$ does not compromise the correctness of the resulting $S_{\mathcal{T}}$ (using either method described before), the resulting $S_{\mathcal{T}}$ will again be more intrusive than if one uses $\varphi_{\mathbb{B}}$, as some precision is lost (in particular for inputs $(x<1)$, which are treated by $\varphi_{\mathbb{B}}^{\prime}$ as $(x<2)$). We now rigorously formalize this intuition.
@@
If $R$ is legitimate, then $R\subseteq\textit{VR}$. If additionally $R$ is a strict covering, then $R=\textit{VR}$. Strict covering implies that all possible moves of the environment are covered, regardless of whether there are moves that a clever environment will never play because these moves leave more power to the system than better, alternative moves. A non-strict covering does not necessarily consider all the possible moves of the environment in the game, but it still considers at least all optimal environment moves. A non-strict covering can be evaluated by checking that with regard to $R$, for all $\overline{x}$, the disjunction of the playable choices (see $\bigwedge_{c\in r}\exists\overline{y}.f_{c}$ of $f_{r}$ in Sec. 2) of its reactions holds.
@@
Proof.
Let $\varphi_{\mathbb{B}}$ created from VR and let $R\subset\textit{VR}$ be a feasible reaction set. This means that there is $r\in\textit{VR}\setminus R$ such that there is an $r^{\prime}\in R$ with $r^{\prime}\subseteq r$. Consider a play of $\varphi_{\mathbb{B}}$ where the environment plays the move corresponding to $r$ at some point. If the move is replaced by $r^{\prime}$, then if the play with $r$ was winning for the environment then the play with $r^{\prime}$ is also winning for the environment. Therefore, if $\varphi_{\mathbb{B}}$ is winning for the environment, there is a strategy for the environment where $r$ is never played. It follows that the $\varphi_{\mathbb{B}}^{\prime}$ that results from $R$ is equi-realizable to the $\varphi_{\mathbb{B}}$ that results from VR. ∎
@@
To obtain an equi-realizable Boolean abstraction, it is not necessary to consider VR, and instead a feasible set is sufficient. The computation of a feasible set is faster, and generates smaller $\varphi_{\mathbb{B}}$ formulae (?). However, there is a price to pay in terms of how permissive the shield is. In practice, it regularly happens that the environment plays moves that are not optimal in the sense that other moves would leave less choice to the system.
@@
Proof.
Let $R$ be a feasible set of reactions. Recall that for every $r\in R$, any $r^{\prime}\neq r$ s.t. $r^{\prime}\in R$ it holds that either $r\subseteq r^{\prime}$ or $r^{\prime}\subseteq r$ or $r$ and $r^{\prime}$ are not comparable. Thus, for any arbitrary $(r,r^{\prime})$ where $r,r^{\prime}\in R$, and $r$ and $r^{\prime}$ are comparable, we can generate a feasible set $R^{\prime}$ where either $r\notin R^{\prime}$ or $r^{\prime}\notin R^{\prime}$; in other words, an exhaustive algorithm can remove all the comparable reactions until the fix-point where all of them are not comparable and the resulting $R^{k}$ is feasible (where $k$ is the number of iterations). Moreover, $R^{\prime\prime}$ is exactly MVR. Also, since synthesis of $C_{\mathbb{B}}$ is decidable, then computing the most intrusive $S_{\mathcal{T}}$ is decidable. ∎
In Ex. 3 the alternative $\varphi_{\mathbb{B}}^{\prime}$ that ignores the (playable) choice $e_{0}^{+}$ is exactly comparing $r_{e_{0}}=\{c_{1},c_{2}\}$ and $r_{e_{0}^{+}}=\{c_{1},c_{2},c_{3}\}$ and constructing a feasible $R^{\prime}$, where $r_{e_{0}^{+}}\notin R^{\prime}$, since $r_{e_{0}}\subset r_{e_{0}^{+}}$ In this case, the fix point is found in only one step and $R^{\prime}$ is an MVR. There are algorithms (Alg. 1 in (?)) that provide a VR set, while others (Alg. 2 and Alg. 3 in (?)) produce feasible reaction sets, which are not necessarily MVR. Our proof of Thm. 4 suggests such (exhaustive) algorithm to provide MVR.
3.3 Further Advantages of Shields in $\mathcal{T}$
@@
In Subsec. 3.1 we observed that, given $v_{\overline{x}}$, the provider component computes an output $v_{\overline{y}}^{\prime}$ from $c$ and $r$ such that $v_{\overline{y}}^{\prime}$ is a model of the formula $\psi=\exists\overline{y}^{\prime}.\big{(}f_{r}(\overline{x})\mathrel{\rightarrow}f_{c}(\overline{y}^{\prime},\overline{x})\big{)}[\overline{x}\leftarrow v_{\overline{x}}]$. Moreover, not only is $\psi$ guaranteed to be satisfiable (by Lemma. 1), but usually has several models. This implies that the engineer can select some $v_{\overline{y}}^{\prime}$ that are preferable over others, depending on different criteria represented by objective functions. Optimization Modulo Theories (OMT) is an extension of SMT that allows model selection based on a given set of constraints (?). We first describe $S_{\mathcal{T}}$ with linear optimization, which is suitable for arithmetic theories.
@@
Recall from Ex. 3 that $x:0$. Also, for simplicity we consider $f_{r}=\top$ and $\psi=\exists y^{\prime}.f_{r}(x)\mathrel{\rightarrow}f_{c}(x,y^{\prime})[x\leftarrow 0]$, where $f_{c}=((x<10)\mathrel{\wedge}\neg(y^{\prime}>9)\mathrel{\wedge}\neg(y^{\prime}\leq x))$. Then, a possible model for $\psi$ is $y^{\prime}:2$. However, any value $y^{\prime}\in[1,9]$ would a valid choice. Therefore, we can augment $\psi$ with objective functions that explore the range $[1,9]$ while maximizing different criteria, which we will denote with $\psi_{f^{+}}$, for some objective function $f^{+}$. For instance, the engineer might be interested in obtaining the greatest $y^{\prime}$ possible in $\psi$, in which case using $f^{+}=\text{max}(y^{\prime})$ would result in $\psi_{f^{+}}$—returning $y^{\prime}:9$. Additionally, we can optimize linear combinations, e.g., $f^{+}=\text{min}(x+y^{\prime})$ in $\psi_{f^{+}}$, which returns $y^{\prime}:1$.
Minimizing distances.
Although $S_{\mathcal{T}}$ with linear optimization allows enriching $S_{\mathcal{T}}$ with better responses, it suffers from a crucial limitation: it cannot solve $\psi_{f^{+}}$, where $f^{+}=\text{min}(|y-y^{\prime}|)$, i.e., the objective function that minimizes the distance between $v_{y}$ by $D$ and $v_{y}^{\prime}$ by $S_{\mathcal{T}}^{\prime}$. This is because these are not linear properties. However, this optimization would be very relevant in the context of shielding, because, without loss of generality, it expresses that $v_{\overline{y}}^{\prime}$ is the safe correction by $S_{\mathcal{T}}$ closest to the unsafe $v_{\overline{y}}$ by $D$. To solve this, we used maximum satisfiability (MaxSMT), which adds soft constraints =$\mathcal{M}=\{\phi_{1},\phi_{2},...\}$ to $\psi$, such that $\psi(\mathcal{M})=\exists\overline{y}^{\prime}.(f_{c}(\overline{x},\overline{y}^{\prime})\prescript{+}{}{\bigwedge}^{|\mathcal{M}|}_{i=0}\phi_{i})[\overline{x}\leftarrow v_{\overline{x}}]$, where $\prescript{+}{}{\bigwedge}$ denotes a soft conjunction, meaning that the right-hand side is satisfied only if possible. To better illustrate this, we use a single variable $y$, although this concept can also be extended to other notions of distance with multiple variables $\overline{y}$ (e.g., Euclidean distance).
Example 5.
Consider again Ex. 3 and let =$\mathcal{M}=\{(y^{\prime}>5)\}$. Then $\psi(\mathcal{M})=\exists y^{\prime}.(f_{c}(x,y^{\prime})\prescript{+}{}{\wedge}(y^{\prime}>5))[x\leftarrow 0]$ does not return $y^{\prime}:2$, but some $y^{\prime}\in[6,9]$. In addition, as stated, we can leverage MaxSMT to express that $v_{y}^{\prime}$ is the safe correction that is the closest to an unsafe $v_{y}$, for which we add $\phi(\overline{x},y,y^{\prime})=\forall z.(f_{c}(\overline{x},z)\mathrel{\rightarrow}(|y^{\prime}-y|<|z-y|))$ to $\mathcal{M}$ in $\psi(\mathcal{M})$, so that $\psi(\mathcal{M})=(f_{c}(\overline{x},y)\wedge(y^{\prime}>5)\wedge\phi(\overline{x},y,y^{\prime}))[\overline{x}\leftarrow v_{\overline{x}},y\leftarrow v_{y}]$. Thus, given a $y:4$ labelled as unsafe, $\psi(\mathcal{M})$ will not return an arbitrary model $y^{\prime}\in[6,9]$, but the concrete $y^{\prime}:6$. This converges in $\mathcal{T}_{\mathbb{Z}}$.
Lemma 4 (Closest element).
In $\mathcal{T}_{\mathbb{Z}}$ the following holds. Assuming $\forall\overline{x}.\exists y.f_{c}(\overline{x},y)$, the following is also valid:
$\forall\overline{x}.z.\exists y.\big{(}f_{c}(\overline{x},y)\mathrel{\wedge}\forall w.[f_{c}(\overline{x},w)\mathrel{\rightarrow}|y-z|\leq|w-z|]).$
In other words, in $\mathcal{T}_{\mathbb{Z}}$ if for all inputs $\overline{x}$ there is an output $y$ such that $\psi$ holds, then there is always a closest value to any provided $z$ that satisfies $\psi$. However, for the theory of real arithmetic $\mathcal{T}_{\mathbb{R}}$ it may not be possible to find a value that satisfies $f_{c}(\overline{x},y^{\prime})$ and is the closest to $v_{y}$. Instead, it is always possible to compute a closest value within a given tolerance constant $\epsilon$ via the addition of the following, soft constraint: $\forall z.(f_{c}(\overline{x},z)\mathrel{\rightarrow}(|y^{\prime}-y|<|z-y|+\epsilon))$.
Lemma 5 (Approximately Closest Element).
In $\mathcal{T}_{\mathbb{R}}$ the following holds. Assuming $\forall\overline{x}.\exists y.f_{c}(\overline{x},y)$, then for every constant $\epsilon>0$, the following is valid:
$\forall\overline{x}.z.\exists y.\big{(}f_{c}(\overline{x},y)\mathrel{\wedge}\forall w.[f_{c}(\overline{x},w)\mathrel{\rightarrow}|y-z|\leq|w-z|+\epsilon]).$
@@
We showed that a $S_{\mathcal{T}}$ can provide an output $v_{\overline{y}}^{\prime}$ that is the safe correction by closest to the unsafe $v_{\overline{y}}$ by $D$. Moreover, previously we showed that we can generate multiple strategies using WR. Therefore, we can optimize $v_{\overline{y}}^{\prime}$ using combinations in the WR; for instance, return $v_{\overline{y}}^{\prime}$ such that distance to $v_{\overline{y}}$ is minimal and $v_{\overline{y}}^{\prime}$ has been chosen among the strategies represented by automata with less than $n$ states (i.e., using bounded synthesis with bound $n$). We illustrate this using our running example.
Example 6.
Consider again $\varphi_{\mathbb{B}}$ and $\varphi_{\mathbb{B}}^{\prime}$ from Ex. 3 and environment input $x:0$. For the sake of the argument, consider that the correct output has to satisfy either $f_{c_{2}}$ (both in $e_{0}$ and $e_{0}^{+}$), or $f_{c_{2}}$ or $f_{c_{3}}$ (where $f_{c_{3}}$ appears only in $e_{0}^{+}$). Let the candidate output $y:11$, which does not correspond to neither $f_{c_{2}}$ or $f_{c_{3}}$, which means that has to be overridden. Now, we want to produce a correction $v_{\overline{y}}^{\prime}$ as close as possible to $v_{\overline{y}}$. If we only consider $e_{0}$, the best $v_{\overline{y}}^{\prime}$ (in the distance sense) that we can produce is $y^{\prime}:0$, because $f_{c_{2}}$ has to hold constraint $(y\leq x$). On the contrary, if we consider $e_{0}$, then the best $v_{\overline{y}}^{\prime}$ we can produce is $y^{\prime}:9$ (a higher output is not possible because of constraint $(y\leq 9$) of $f_{c_{3}}$).
@@
4 Evaluation and Practical Implications
4.1 Completeness in Fully Observable Domains
Case Study: GridWorld. We demonstrate that in scenarios that are fully observable, $S_{\mathcal{T}}$ can be complete in the sense that it is possible to fully guarantee safety. Gridworld (GW) is a standard DRL benchmark (?), in which an agent navigates on a two-dimensional grid (see leftmost side of Fig. 2) and its goal is to reach a target destination while adhering to some additional requirements. Typically, GW has a discrete state space representing the coordinates of the cells. However, to show the capabilities of our shielding approach, we modify the scenario to obtain a continuous state space by adding Gaussian noise to the agent’s position, resulting in inputs in the domain of $\mathcal{T}_{\mathbb{R}}$. We note that this also makes the problem significantly harder. One main advantage of GW is that it is a fully observable system, i.e., we can manually encode a complete set of requirements that, if respected, guarantee that $D\cdot S_{\mathcal{T}}$ never violates a high-level behaviour, e.g., never collide. In other words, $S_{\mathcal{T}}$ can only enforce $D\cdot S_{\mathcal{T}}$ based on $\varphi_{\mathcal{T}}$, whereas successfully avoiding collisions depends on whether the properties captured by $\varphi_{\mathcal{T}}$ cover all possible traces. This is possible in GW and other fully observable systems. We chose the natural safety specification — do not collide at the next time-step, and we demonstrate how our shield can guarantee that this specification is never violated.
Figure 2: GridWorld, left to right: A scheme of the benchmark arena, and the average success rate and collision rate obtained during the DRL training process.
Results. We trained a pool of $500$ models, differing in the random seed used to generate them. Next, we retained only the models that demonstrated very high performance, i.e., reached our cutoff threshold of a $98\%$ success rate, with successful trajectories in over $100$ consecutive trials (see center images in Fig. 2). We also point out that even after extensive training, there was not a single model that achieved a $0\%$ collision rate (see rightmost image in Fig. 2). We note that this is in line with previous research on DRL safety, which demonstrated that even state-of-the-art agents may indeed violate simple safety requirements (???) — further motivating the need for shielding even in this, relatively simple, task. Next, we showcase that $S_{\mathcal{T}}$ can efficiently eliminate collisions via the usage of safety properties that cause the agent to collide in the next time-step, and overriding these, unwanted actions. For example, if the agent is near an obstacle on the left   [^3] Represented with the continuous input $\overline{x}=\{x_{0},x_{1},...,x_{7}\}$, where $x_{0}\in[0.05,0.15]$, $x_{1}\in[0.05,0.15]$, …, $x_{7}\in[0.05,0.15]$. then it should not select the next action to be LEFT. We now analyze the runtime impact of $S_{\mathcal{T}}$ on real executions of $D\cdot S_{\mathcal{T}}$. Tab. 1 present the first $10$ results, where we can see that our proposed solution successfully eliminates all possible collisions, resulting in safer behavior. Nevertheless, we point out that the increased safety did not compromise the general performance of the agent, as can be seen by the unaltered success rate.
Seed Without Shield With Shield Overhead
Succ. Coll. Succ. Coll.
0 $89\%$ $3\%$ $89\%$ $0\%$ $219.1\%$
1 $92\%$ $1\%$ $93\%$ $0\%$ $182.4\%$
2 $91\%$ $2\%$ $92\%$ $0\%$ $204.6\%$
3 $91\%$ $3\%$ $91\%$ $0\%$ $240.9\%$
4 $92\%$ $2\%$ $92\%$ $0\%$ $206.5\%$
5 $89\%$ $1\%$ $90\%$ $0\%$ $213.1\%$
6 $92\%$ $2\%$ $92\%$ $0\%$ $212.1\%$
7 $86\%$ $6\%$ $91\%$ $0\%$ $250.9\%$
8 $92\%$ $1\%$ $92\%$ $0\%$ $226.2\%$
9 $89\%$ $3\%$ $91\%$ $0\%$ $220.0\%$
Table 1: A performance comparison for the agent with and without the safety shield. The values are obtained over $100$ episodes for each of the seeds.
Although $D\cdot S_{\mathcal{T}}$ completely avoids collisions, this comes at a price: $D\cdot S_{\mathcal{T}}$ checks in each time-step whether $(v_{\overline{x}},v_{\overline{y}})$ is correct with respect to $\varphi_{\mathcal{T}}$ (and overrides it if not). This, in turn, introduces a computational overhead in the execution. In the case of GW, we note that the overhead is significant (around 200%), which suggests that the choice of using a more permissive $S_{\mathcal{T}}$ (see Subsec. 3.2), might be preferable in domains where runtime optimization is essential. Note that success ratios do not necessarily increase to $100\%$ when shielded, as collision avoidance may cause some loops instead of collisions. Nevertheless success never decreases.
4.2 Temporal Dynamics

### v2

@@ Introduction
Note that if the environment generates an input $v_{\overline{x}}$ (with corresponding $v_{\overline{e}}$) and $D$ responds with an output $v_{\overline{y}}$ such that $\textsf{getchoice}(v_{\overline{y}})=v_{\overline{s}}$, if $W$ permits $(v_{\overline{e}},v_{\overline{s}})$ then the suggested $v_{\overline{y}}$ remains. Otherwise, we use a component find which computes an alternative $v_{\overline{s}}^{\prime}$ such that $(v_{\overline{e}},v_{\overline{s}}^{\prime})$ is in $W$ and a proper value $v_{\overline{y}}$ that corresponds to $v_{\overline{s}}$ is output. Also, note that find can always identify a value $v_{\overline{s}}^{\prime}$ from any $q\in Q_{\textit{now}}$ such that there is a successor $(q^{\prime},v_{\overline{s}}^{\prime})\in\delta(q,v_{\overline{e}})$ by the definition of WR. Given two $v_{\overline{s}}$ and $v_{\overline{s}}^{\prime}$ such that $(v_{\overline{e}},v_{\overline{s}})\in W$ and $(v_{\overline{e}},v_{\overline{s}}^{\prime})\in W$, the criteria (e.g., choice-logic (Bernreiter, Maly, and Woltran 2021)) to choose one is not relevant for this paper and we leave it for future work. Finally, given $v_{\overline{x}}$, the $W$ cannot distinguish two values $v_{\overline{y}}$ and $v_{\overline{y}}^{\prime}$ if $\textsf{getchoice}(v_{\overline{x}},v_{\overline{y}})=\textsf{getchoice}(v_{\overline{x}},v_{\overline{y}}^{\prime})$.
@@
Let $\varphi_{\mathcal{T}}$ be a safety specification and $\varphi_{\mathbb{B}}$ an abstraction of $\varphi_{\mathcal{T}}$. Let $S_{\mathcal{T}}$ be synthesized using the winning region $W$ of $\varphi_{\mathbb{B}}$. Then, $D\cdot S_{\mathcal{T}}$ satisfies $\varphi_{\mathcal{T}}$, for any external controller $D$.
@@
Lemma 3 essentially states that a WR controller is less intrusive than one based on a specific Boolean controller $C_{\mathbb{B}}$. However, in practice, a WR is harder to interpret and slower to compute than $C_{\mathbb{B}}$. Note that we presented $S_{\mathcal{T}}$ based on $C_{\mathcal{T}}$ and $S_{\mathcal{T}}$ based on WR, but for classic LTL shielding only the latter case makes sense. This is because in classic LTL using shield $S$ based on $C_{\mathbb{B}}$ is equivalent to forcing $D$ to coincide with $S$ on all time-steps, so $S$ could be directly used ignoring $D$. However, in $\textup{LTL}_{\mathcal{T}}$ there are potentially many different $v_{\overline{y}}$ by $D$ that correspond to a given Boolean $v_{\overline{s}}^{\prime}$ by $C_{\mathbb{B}}$, so even when this architecture forces to mimick $C_{\mathbb{B}}$ in the Boolean domain, it is reasonable to use an external source $D$ of rich value output candidates. Thus, not only we present an extension of shielding to $\textup{LTL}_{\mathcal{T}}$ (with the WR approach), but also this is the first work that is capable to synthetise shields leveraging all the power of classic synthesis (with the $C_{\mathcal{T}}$ approach): e.g., we can use bounded synthesis (Schewe and Finkbeiner 2007) instead of computing a WR.
Permissive and Intrusive Shields
@@
Computing a Boolean Abstraction boils down to calculating the set VR of valid reactions. In order to speed up the calculation, optimizations are sometimes used (see (Rodriguez and Sánchez 2023)). For instance, Ex. 1 shows a formula $\varphi_{\mathcal{T}}$ that can be Booleanized into a $\varphi_{\mathbb{B}}$ that contains three environment variables ($e_{0}$, $e_{0}^{+}$ and $e_{1}$) that correspond to the three valid reactions. A closer inspection reveals that the valid reaction associated to $e_{0}^{+}$ (which represents $(x<1)$) is “subsumed” by the reaction associated to $e_{0}$ (which represents $(x<2)$), which is witnessed by $e_{0}$ leaving to the system a strictly smaller set of choices than $e_{0}^{+}$. In the game theoretic sense, an environment that plays $e_{0}^{+}$ is leaving the system strictly more options, so a resulting Booleanizaed formula $\varphi_{\mathbb{B}}^{\prime}$ that simply ignores $e_{0}^{+}$ only limits the environment from playing sub-optimal moves, and $\varphi_{\mathbb{B}}^{\prime}$ is also equi-realizable to $\varphi_{\mathcal{T}}$. From the equi-realizability point of view, one can neglect valid reactions that provide strictly more choices than other valid reactions. However, even though using $\varphi_{\mathbb{B}}^{\prime}$ does not compromise the correctness of the resulting $S_{\mathcal{T}}$ (using either method described before), the resulting $S_{\mathcal{T}}$ will again be more intrusive than if one uses $\varphi_{\mathbb{B}}$, as some precision is lost (in particular for inputs $(x<1)$, which are treated by $\varphi_{\mathbb{B}}^{\prime}$ as $(x<2)$). We now rigorously formalize this intuition.
@@
If $R$ is legitimate, then $R\subseteq\textit{VR}$. If additionally $R$ is a strict covering, then $R=\textit{VR}$. Strict covering implies that all possible moves of the environment are covered, regardless of whether there are moves that a clever environment will never play because these moves leave more power to the system than better, alternative moves. A non-strict covering does not necessarily consider all the possible moves of the environment in the game, but it still considers at least all optimal environment moves. A non-strict covering can be evaluated by checking that with regard to $R$, for all $\overline{x}$, the disjunction of the playable choices (see $\bigwedge_{c\in r}\exists\overline{y}.f_{c}$ of $f_{r}$ in Preliminaries) of its reactions holds.
@@
To obtain an equi-realizable Boolean abstraction, it is not necessary to consider VR, and instead a feasible set is sufficient. The computation of a feasible set is faster, and generates smaller $\varphi_{\mathbb{B}}$ formulae (Rodriguez and Sánchez 2023). However, there is a price to pay in terms of how permissive the shield is. In practice, it regularly happens that the environment plays moves that are not optimal in the sense that other moves would leave less choice to the system.
@@
In Ex. 3 the alternative $\varphi_{\mathbb{B}}^{\prime}$ that ignores the (playable) choice $e_{0}^{+}$ is exactly comparing $r_{e_{0}}=\{c_{1},c_{2}\}$ and $r_{e_{0}^{+}}=\{c_{1},c_{2},c_{3}\}$ and constructing a feasible $R^{\prime}$, where $r_{e_{0}^{+}}\notin R^{\prime}$, since $r_{e_{0}}\subset r_{e_{0}^{+}}$ In this case, $R^{\prime}$ is an MVR.
Optimizations in $\textup{LTL}_{\mathcal{T}}$ Shields
@@
We know that, given $v_{\overline{x}}$, the provider component computes an output $v_{\overline{y}}^{\prime}$ from $c$ and $r$ such that $v_{\overline{y}}^{\prime}$ is a model of the formula $\psi=\exists\overline{y}^{\prime}.\big{(}f_{r}(\overline{x})\mathrel{\rightarrow}f_{c}(\overline{y}^{\prime},\overline{x})\big{)}[\overline{x}\leftarrow v_{\overline{x}}]$. Moreover, not only is $\psi$ guaranteed to be satisfiable (by Lemma. 1), but usually has several models. This implies that the engineer can select some $v_{\overline{y}}^{\prime}$ that are preferable over others, depending on different criteria represented by objective functions, such as $\psi_{f^{+}}$, where $f^{+}=\text{min}(|y-y^{\prime}|)$, i.e., the objective function that minimizes the distance between $v_{y}$ by $D$ and $v_{y}^{\prime}$ by $S_{\mathcal{T}}^{\prime}$. This is because these are not linear properties. However, this optimization would be very relevant in the context of shielding, because, without loss of generality, it expresses that $v_{\overline{y}}^{\prime}$ is the safe correction by $S_{\mathcal{T}}$ closest to the unsafe $v_{\overline{y}}$ by $D$. To solve this, we used maximum satisfiability, which adds soft constraints =$\mathcal{M}=\{\phi_{1},\phi_{2},...\}$ to $\psi$, such that $\psi(\mathcal{M})=\exists\overline{y}^{\prime}.(f_{c}(\overline{x},\overline{y}^{\prime})^{+}{}\bigwedge^{|\mathcal{M}|}_{i=0}\phi_{i})[\overline{x}\leftarrow v_{\overline{x}}]$, where ${}^{+}{}\bigwedge$ denotes a soft conjunction, meaning that the right-hand side is satisfied only if possible. To better illustrate this, we use a single variable $y$, although this concept can also be extended to other notions of distance with multiple variables $\overline{y}$ (e.g., Euclidean distance) In our extended paper (Rodriguez et al. 2024) we include a formal discussion on how to use optimizations depending on $\mathcal{T}$..
@@
Consider again Ex. 3 and let =$\mathcal{M}=\{(y^{\prime}>5)\}$. Then $\psi(\mathcal{M})=\exists y^{\prime}.(f_{c}(x,y^{\prime})^{+}{}\wedge(y^{\prime}>5))[x\leftarrow 0]$ does not return $y^{\prime}:2$, but some $y^{\prime}\in[6,9]$. In addition, we use maximum satisfiability to express that $v_{y}^{\prime}$ is the safe correction that is the closest to an unsafe $v_{y}$, for which we add $\phi(\overline{x},y,y^{\prime})=\forall z.(f_{c}(\overline{x},z)\mathrel{\rightarrow}(|y^{\prime}-y|<|z-y|))$ to $\mathcal{M}$ in $\psi(\mathcal{M})$, so that $\psi(\mathcal{M})=(f_{c}(\overline{x},y)\wedge(y^{\prime}>5)\wedge\phi(\overline{x},y,y^{\prime}))[\overline{x}\leftarrow v_{\overline{x}},y\leftarrow v_{y}]$. Thus, given a $y:4$ labelled as unsafe, $\psi(\mathcal{M})$ will not return an arbitrary model $y^{\prime}\in[6,9]$, but the concrete $y^{\prime}:6$. This converges in $\mathcal{T}_{\mathbb{Z}}$.
@@
We showed that a $S_{\mathcal{T}}$ can provide an output $v_{\overline{y}}^{\prime}$ that is the safe correction by closest to the unsafe $v_{\overline{y}}$ by $D$. Moreover, previously we showed that we can generate multiple strategies using WR. Therefore, we can optimize $v_{\overline{y}}^{\prime}$ using combinations in the WR; for instance, return $v_{\overline{y}}^{\prime}$ such that distance to $v_{\overline{y}}$ is minimal and $v_{\overline{y}}^{\prime}$ has been chosen among the strategies represented by automata with less than $n$ states (i.e., using bounded synthesis with bound $n$). We illustrate this in (Rodriguez et al. 2024).
@@
Related Work and Conclusion
Related Work.
Classic shielding approaches (Bloem et al. 2015; Könighofer et al. 2017; Alshiekh et al. 2018; Avni et al. 2019) focus on properties expressed in Boolean LTL, and are incompatible for systems with richer-data domains: i.e., they need explicit manual discretization of the requirements. In the other hand, we have a sound procedure that directly takes $\textup{LTL}_{\mathcal{T}}$ specifications (via (Rodriguez and Sánchez 2024)) which we adapted to shielding. This is fundamentally new, and also that we can also optimize outputs with respect to erroneous candidates.
Competing methods